OpenStudy (anonymous):

A street light is mounted at the top of a 14 ft tall pole. A woman 6 ft tall walks away from the pole with a speed of 6 ft/sec along a straight path. How fast is the tip of her shadow moving when she is 45 ft from the base of the pole?

OpenStudy (anonymous):

use similar triangles. create two similar triangles based on the information tehn take the derivative

OpenStudy (anonymous):

i dont understand how to do that

OpenStudy (anonymous):

check out this website and go to example 2, it has a similar question with different numbers =) http://www.math.wfu.edu/tutorials/Math111/RelatedRates.pdf

OpenStudy (bahrom7893):

okay carra i will solve it, im done with my hw

OpenStudy (bahrom7893):

For the answer try 9/2 (ft/sec) , let me know whether it is correct. If it is, I will post the solution.

OpenStudy (anonymous):

no not the answer

OpenStudy (bahrom7893):

try 4.5

OpenStudy (anonymous):

no

OpenStudy (bahrom7893):

k let me see if i did anythinig wrong, brb

OpenStudy (bahrom7893):

can you double check the question?

OpenStudy (anonymous):

question is correct

OpenStudy (bahrom7893):

okay, let me check my work again

OpenStudy (bahrom7893):

can you try 4/3 or 1.3

OpenStudy (anonymous):

no i sorry

OpenStudy (anonymous):

hey

OpenStudy (anonymous):

im on the case

OpenStudy (anonymous):

do you mind, i will use twiddla

OpenStudy (anonymous):

http://www.twiddla.com/492533

OpenStudy (anonymous):

click on that

OpenStudy (bahrom7893):

Okay so cantorset, this is what im doing: y/6 = (y+x)/14 => look at that in terms of similar triangles. Cross multiply and simplify: 14y = 6x + 6y 14(dy/dt)=6(dx/dt) + 6(dy/dt) 8 d(y/dt) = 6 (dx/dt) 8(dy/dt) = 6 * 6 8(dy/dt) = 36, solve for dy/dt = 9/2. What am I doing wrong?

OpenStudy (anonymous):

its a whiteboard

OpenStudy (bahrom7893):

carra try 10.5

OpenStudy (bahrom7893):

uploading the screen

OpenStudy (bahrom7893):

http://i53.tinypic.com/o6i8b7.jpg there's the screen

OpenStudy (bahrom7893):

i gtg now, bbl